The consumer has a budget of $M for the consumption of good X and Y .The price of good X is $Px and the price of good Y is $py.The consumer has a utility function given by U(X,Y)=xy.

Find the optimal consumption of the individual and the utility Obtained.

Make a graph that illustrates the solution to the problem.

Why do you use paramerters.


1
Expert's answer
2021-10-15T10:20:31-0400

The optimal consumption of the individual occurs when:

"MUx\/MUy = Px\/Py" and "Px\u00d7x + Py\u00d7y = M."

"MUx = U'(X) = y,"

"MUy = U'(y) = x,"

y/x = Px/Py,

y = Px×x/Py or Px×x = Py×y,

2Px×x = 2Py×y = M,

y = 0.5M/Py,

x = 0.5M/Px.